# VideoLLaMA 2 Advancing Spatial-Temporal Modeling and Audio Understanding in Video-LLMs

### v1

@@ 4.1.1 Evaluation Benchmarks
For the video captioning task, we perform experiments on the newly introduced Multi-Source Video Caption (MSVC) benchmark. MSVC is introduced to address limitations in existing video caption benchmarks, MSVC samples 500 videos with human-annotated captions from MSVD (Chen & Dolan, 2011), MSRVTT (Xu et al., 2016), and VATEX (Wang et al., 2019), ensuring diverse scenarios and domains. Traditional evaluation metrics rely on exact match statistics between generated and ground truth captions, which are limited in capturing the richness of video content. Thus, we use a ChatGPT-assisted evaluation similar to Maaz et al. (2023b). Both generated and human-annotated captions333Each video in the MSVC benchmark is annotated with multiple human-written captions, covering different aspects of the video. This comprehensive annotation ensures a robust and thorough evaluation of Video-LLMs. are evaluated by GPT-3.5-turbo (0613) for Correctness of Information and Detailed Orientation.
@@
| Gemini 1.5 Flash (Google, 2024) | - | - | - | - | - | 3.46♠ | 3.24♠ |
| Gemini 1.5 Pro (Google, 2024) | - | 63.2 | - | - | 75.7♢ | 3.67♠ | 3.52♠ |
| GPT4-V (OpenAI, 2023b) | - | 55.6 | - | 43.7♢ | 60.7♢ | 2.70♠ | 2.76♠ |
| GPT4-O (OpenAI, 2024) | - | 72.2 | - | - | 66.2♢ | - | - |
@@
| LLaMA-VID (7B) | 1 fps | 38.5♠ | 44.6♠ | 41.9♠ | 25.9♠ | 1.84♠ | 2.11♠ |
| Video-LLaVA (7B) | 8 | 38.4♠ | 44.3♠ | 41.0♠ | 40.4♢ | 1.85♠ | 2.05♠ |
| VideoChat2 (7B) | 16 | 42.2♠ | 47.3♠ | 51.1 | 33.7♢ | 2.01♠ | 2.10♠ |
@@
| VideoLLaMA 2 (7B) | 16 | 51.7 | 51.4 | 54.6 | 46.6 | 2.53 | 2.59 |
| VideoLLaMA 2 (8x7B) | 8 | 53.3 | 52.2 | 53.9 | 48.4 | 2.53 | 2.56 |
@@ 4.1.2 Baselines
The proprietary models selected for comparison are state-of-the-art multi-modal systems developed by leading companies. These models include Gemini 1.0 Series (Google, 2023), Gemini 1.5 Series (Google, 2024), GPT4-V (OpenAI, 2023b), GPT4-O (OpenAI, 2024), and Reka Series (Reka, 2024). These models represent the cutting-edge proprietary multi-modal understanding technologies and serve as benchmarks to gauge the performance of VideoLLaMA 2.
@@
We also include several prominent open-source models to provide a broader context for our evaluations. Specifically, we compare our VideoLLaMA 2 with VistaLLaMA (Ma et al., 2023), ChatUniVi (Jin et al., 2023b), LLaMA-VID (Li et al., 2023d), Video-LLaVA (Lin et al., 2023a), VideoChat2 (Li et al., 2023c), and LLaVA-NeXT-Video Series444For a fair comparison, we do not include the preference-optimized LLaVA-NeXT-Video as baseline. However, we still provide the full comparison results between VideoLLaMA 2 and LLaVA-NeXT-Video-DPO in Appendix B for the reference of readers. (Liu et al., 2024b). These open-source models are crucial for evaluating the performance of VideoLLaMA 2 within the context of accessible and reproducible research.
@@ 4.2.1 Evaluation Benchmarks
For open-ended audio-video question answering, we adopt AVSSD (Chen et al., 2020) (15341 samples), AVSD (Alamri et al., 2019) (18630 samples) and Music-AVQA (Li et al., 2022a) (9129 samples) as benchmarks. For AVSSD and AVSD, we employ a GPT-assisted evaluation as the same as the protocols in OE-VQA. We randomly sample one-eighth of the dataset as a subset for evaluation since the full dataset is too large for ChatGPT inference.
@@ 4.2.2 Baselines
To assess the audio understanding capabilities of VideoLLaMA 2-7B, we utilize Qwen-Audio (Chu et al., 2023) as a primary benchmark. Qwen-Audio is renowned for its robust performance across various audio understanding tasks, providing a strong comparison point for our model.
@@
Expanding the evaluation to audio-visual integration, VideoLLaMA 2 is benchmarked against advanced models that specialize in handling multimodal data. This includes comparisons with PandaGPT (Su et al., 2023), Video-LLaMA (Zhang et al., 2023) and MacawLLM (Lyu et al., 2023), which are adept at understanding complex multimodal scenes. Further, we examine newer integrative models like AV-LLM (Shu et al., 2023) and AVicuna (Tang et al., 2024), which utilize high-quality audio-visual training datasets to enhance their multimodal understanding capabilities. For the AVSD and AVQA-Music datasets, these baseline models are evaluated in a zero-shot manner.
@@ 4.2.3 Main Results
| Method | # Hours | Clotho-AQA | TUT2017 |
| Qwen-Audio (7B) | 137k | 57.90 | 64.90 |
@@
| Method | # Pairs | MUSIC-QA | AVSD | AVSSD |
@@
The experimental findings presented in Table 7 delineate the compelling advantages of VideoLLaMA 2 (7B) in the field of audio-only question answering (AQA).
Using the widely recognized Qwen-Audio (Chu et al., 2023) as a benchmark, VideoLLaMA 2-7B demonstrates remarkable performance, particularly within the Clotho-AQA dataset. Despite being trained on a significantly smaller dataset—just 4k hours of audio data—VideoLLaMA 2-7B surpasses Qwen-Audio-7B by a notable margin of 2.7%. In the context of the TUT2017 dataset, which features multiple-choice audio question answering, VideoLLaMA 2-7B achieves performance comparable to that of Qwen-Audio. These outcomes underscore the sophisticated capabilities of VideoLLaMA 2-7B, showcasing its ability to effectively understand and interpret audio data with much less initial training compared to established models.
@@
The performance results summarized in Table 8 highlight VideoLLaMA 2’s significant advancements compared to open-source models. This analysis underscores VideoLLaMA 2’s capabilities across various AVQA benchmarks, emphasizing its superior performance.
In the MUSIC-QA benchmark, VideoLLaMA 2-7B not only surpasses other models but does so with a significant margin, registering an impressive score of 70.6%. This score indicates a robust ability to synthesize and analyze complex musical and audio cues, which is central to the MUSIC-QA benchmark. The model’s performance, far outreaching AVicuna (7B) at 49.6%, underscores its advanced capabilities in handling intricate audio-visual interactions. In the AVSD benchmark, which focuses on audio-visual scene description involving dialogue understanding, VideoLLaMA 2-7B scores 50.9%. Although this score does not lead among the competition, with AVicuna (7B) achieving a slightly higher score of 53.1%, it remains competitive. The AVSSD benchmark, which necessitates a deep understanding of both visual scenes and their auditory elements, sees VideoLLaMA 2-7B leading with a score of 62.9%. This performance, much higher than AV-LLM’s 47.6%, showcases VideoLLaMA 2-7B’s exceptional ability to interpret and synthesize information across both modalities effectively. It indicates a profound understanding of dynamic interactions and the contextual synthesis necessary in AVSSD tasks.
@@ 6 Related Works
The vision-language adapters of Image-LLMs, such as Cross-Attention (Alayrac et al., 2022; Bai et al., 2023), Q-Former (Li et al., 2023a; Zhu et al., 2023b), Linear Projection (Liu et al., 2023b; Chen et al., 2023a) and Dynamic Visual Tokenizer (Jin et al., 2023c), have been widely adopted in popular Video-LLMs (Zhang et al., 2023; Maaz et al., 2023b; Lin et al., 2023a; Li et al., 2023c; Jin et al., 2024; Ataallah et al., 2024). Despite their reasonable results, these designs are still insufficient for Video-LLMs because they completely ignore temporal aggregation and leave everything of temporal modeling to language decoder, which is neither computationally efficient (i.e., LLMs need to process much more tokens) nor effective on video understanding tasks, as observed in our preliminary experiments (See Table 1).
@@
Furthermore, expanding on the core capabilities of Video-LLMs, recent advancements have focused on integrating audio streams, a rich source of contextual cues that are vital for a complete video scene understanding. Models such as PandaGPT (Su et al., 2023) and Video-LLaMA (Zhang et al., 2023) utilize pre-trained systems like ImageBind (Girdhar et al., 2023) for a universal modality alignment, while MacawLLM (Lyu et al., 2023) optimizes separate branches for visual and audio data using diverse instructional sets. Other models like AV-LLM (Shu et al., 2023) and AVicuna (Tang et al., 2024) leverage integrated audio-visual training data to further refine their understanding of complex multimodal content. These developments represent significant strides toward creating more versatile and capable LLMs that can navigate and interpret intricate temporal dynamics with synchronous audio streams effectively.
@@ 7 Conclusion
We present the VideoLLaMA 2 series, a set of generalist Video Large Language Models designed to advance multimodal research in the arena of video-language modeling. By incorporating a meticulously designed Spatial-Temporal Convolution (STC) connector and a jointly trained Audio Branch, VideoLLaMA 2 consistently improves multimodal comprehension across various video and audio-oriented tasks. It outperforms the open-source models of similar size across multiple benchmarks and, in several aspects, achieves performance levels comparable to proprietary models, demonstrating impressive capabilities in modeling temporal dynamics with synchronous audio streams. Furthermore, as a foundational model, VideoLLaMA 2 can be further developed to benefit various more specialized but challenging problems, like long video understanding (Ren et al., 2023; Song et al., 2023; Wang et al., 2024c), video agent (Lin et al., 2023b; Fan et al., 2024; Wang et al., 2024a; He et al., 2024), autonomous driving (Xu et al., 2023; Shao et al., 2024), motion understanding (Wu et al., 2024; Chen et al., 2024a), and robotic manipulation (Liu et al., 2024d).

### v3

@@ 5.1.1 Evaluation Benchmarks
For the video captioning task, we perform experiments on the newly introduced Multi-Source Video Caption (MSVC) benchmark. MSVC is introduced to address limitations in existing video caption benchmarks, MSVC samples 500 videos with human-annotated captions from MSVD (Chen & Dolan, 2011), MSRVTT (Xu et al., 2016), and VATEX (Wang et al., 2019), ensuring diverse scenarios and domains. Traditional evaluation metrics rely on exact match statistics between generated and ground truth captions, which are limited in capturing the richness of video content. Thus, we use a ChatGPT-assisted evaluation similar to Maaz et al. (2023b). Both generated and human-annotated captions101010Each video in the MSVC benchmark is annotated with multiple human-written captions, covering different aspects of the video. This comprehensive annotation ensures a robust and thorough evaluation of Video-LLMs. are evaluated by GPT-3.5-turbo (0613) for Correctness of Information and Detailed Orientation 111111We also include evaluation results on several other popular video benchmarks (Zhou et al., 2024; Li et al., 2024c; Fang et al., 2024; Wang et al., 2024a) in Appendix C..
@@
| Gemini 1.5 Flash (Google, 2024) | - | - | - | - | 70.3/75.0♢ | 3.46♠ | 3.24♠ |
| Gemini 1.5 Pro (Google, 2024) | - | 63.2 | - | - | 75.0/81.3♢ | 3.67♠ | 3.52♠ |
| GPT4-V (OpenAI, 2023b) | - | 55.6 | - | 43.7♢ | 59.9/63.3♢ | 2.70♠ | 2.76♠ |
| GPT4-O (OpenAI, 2024) | - | 72.2 | - | - | 71.9/77.2♢ | - | - |
@@
| LLaMA-VID (7B) | 1 fps | 38.5♠ | 44.6♠ | 41.9♠ | 25.9/-♠ | 1.84♠ | 2.11♠ |
| Video-LLaVA (7B) | 8 | 38.4♠ | 44.3♠ | 41.0♠ | 39.9/41.6♢ | 1.85♠ | 2.05♠ |
| VideoChat2 (7B) | 16 | 42.2♠ | 47.3♠ | 51.1 | 33.7/-♢ | 2.01♠ | 2.10♠ |
| LLaVA-NeXT-Video (7B) | 32 | 43.9♠ | 48.8♠ | 46.5♠ | - | 2.40♠ | 2.52♠ |
| LLaVA-NeXT-Video (32B) | 32 | 60.9 | - | - | 60.2/63.0 | - | - |
| PLLaVA (34B) | 16 | - | - | 58.1 | - | - | - |
| VILA 1.5 (34B) | - | 58.0♠ | - | - | 62.3/64.1♢ | - | - |
| LLaVA-OneVision (72B) | 32 | 62.0 | - | 59.5 | 66.3/69.6 | - | - |
@@
| VideoLLaMA2 (7B) | 16 | 51.7 | 51.4 | 54.6 | 47.9/50.3 | 2.53 | 2.59 |
| VideoLLaMA2 (8x7B) | 8 | 53.3 | 52.2 | 53.9 | 47.9/49.7 | 2.53 | 2.56 |
| VideoLLaMA2 (72B) | 16 | 63.9 | 57.5 | 62.0 | 61.4/63.1 | 2.71 | 2.67 |
| VideoLLaMA2.1 (7B) | 16 | 53.1 | 54.9 | 57.3 | 54.9/56.4 | 2.87 | 2.81 |
@@ 5.1.2 Baselines
The proprietary models selected for comparison are state-of-the-art multi-modal systems developed by leading companies. These models include Gemini 1.0 Series (Google, 2023), Gemini 1.5 Series (Google, 2024), GPT4-V (OpenAI, 2023b), GPT4-O (OpenAI, 2024), Reka Series (Reka, 2024), and Pegasus-1 (Jung et al., 2024). These models represent the cutting-edge proprietary multi-modal understanding technologies and serve as benchmarks to gauge the performance of VideoLLaMA 2.
@@
We also include several prominent open-source models to provide a broader context for our evaluations. Specifically, we compare our VideoLLaMA 2 with VistaLLaMA (Ma et al., 2023), ChatUniVi (Jin et al., 2023b), LLaMA-VID (Li et al., 2023d), Video-LLaVA (Lin et al., 2023a), VideoChat2 (Li et al., 2023c), LLaVA-NeXT-Video Series121212For a fair comparison, we do not include the preference-optimized LLaVA-NeXT-Video as baseline. However, we still provide the full comparison results between VideoLLaMA 2 and LLaVA-NeXT-Video-DPO in Appendix B for the reference of readers. (Liu et al., 2024b), VILA 1.5 (Lin et al., 2024), PLLaVA (Xu et al., 2024a), and LLaVA-OneVision (Li et al., 2024a). These open-source models are crucial for evaluating the performance of VideoLLaMA 2 within the context of accessible and reproducible research.
@@ 5.1.4 Main Results
| LLaVA-NeXT-Video (32B) | 32 | - | 54.3/- | - | - | - | - | - |
| PLLaVA (34B) | 16 | - | 60.9/- | 3.60 | 3.20 | 3.90 | 2.67 | 3.25 |
| LLaVA-OneVision (72B) | 32 | - | 62.3/- | - | - | - | - | - |
@@ 5.2.1 Evaluation Benchmarks
For open-ended audio-video question answering, we adopt VGGSound (Chen et al., 2020) (15341 samples), AVSD (Alamri et al., 2019) (18630 samples) and Music-AVQA (Li et al., 2022a) (9129 samples) as benchmarks. For VGGSound and AVSD, we employ a GPT-assisted evaluation as the same as the protocols in OE-VQA.
@@ 5.2.2 Baselines
To assess the audio understanding capabilities of VideoLLaMA 2-7B, we utilize Qwen-Audio (Chu et al., 2023) and Qwen2-Audio (Chu et al., 2024) as the benchmarks. Qwen-Audio and Qwen2-Audio are renowned for its robust performance across various audio understanding tasks, providing a strong comparison point for our model.
@@
Expanding the evaluation to audio-visual integration, VideoLLaMA 2 is benchmarked against advanced models that specialize in handling multimodal data. This includes comparisons with PandaGPT (Su et al., 2023), MacawLLM (Lyu et al., 2023), Video-LLaMA (Zhang et al., 2023), and X-InstructBLIP (Panagopoulou et al., 2023), which are adept at understanding complex multimodal scenes. Further, we examine newer integrative models like AV-LLM (Shu et al., 2023), OneLLM (Han et al., 2024), CREMA (Yu et al., 2024), and AVicuna (Tang et al., 2024), which utilize high-quality audio-visual training datasets to enhance their multimodal understanding capabilities.
@@ 5.2.3 Main Results
| Method | # Hours | Clotho-AQA | TUT2017 | VocalSound |
| Qwen-Audio (7B) | 137k | 57.90 | 64.90 | 92.89 |
| Qwen2-Audio (7B) | 520k | - | - | 93.92 |
@@
| VideoLLaMA2.1-AV (7B) | 5k | 71.02 | 77.28 | 92.40 |
| Method | # Pairs | MUSIC-QA | AVSD | VGGSound |
@@
The experimental findings presented in Table 7 delineate the compelling advantages of VideoLLaMA 2-AV (7B) and VideoLLaMA 2-AV (7B) on audio-only question answering (AQA).
Using the widely recognized Qwen-Audio and Qwen2-Audio as benchmarks, both VideoLLaMA 2-A (7B) and VideoLLaMA 2-AV (7B) demonstrate exceptional performance across multiple datasets, showcasing remarkable efficiency and advanced learning capabilities despite being trained on significantly fewer hours of data. Specifically, VideoLLaMA 2-A (7B), with only 4k hours of audio data, outperforms Qwen-Audio (137k hours) across most benchmarks. It achieves 68.90% on Clotho-AQA, a substantial improvement over Qwen-Audio’s 57.90%, and records 75.19% on TUT2017, surpassing Qwen-Audio’s 64.90% by 10.29%. This highlights its robust understanding and interpretative capabilities in complex audio question-answering tasks. Meanwhile, in the VocalSound benchmark, where high accuracy is essential, VideoLLaMA 2-A (7B) achieves 92.73%, closely matching Qwen2-Audio (7B)’s 93.92% despite being trained on far fewer hours (4k vs. 520k).
Furthermore, VideoLLaMA 2-AV (7B), leveraging about 5k hours of multimodal (audio-visual) data, builds on the strengths of VideoLLaMA 2-A and achieves even higher performance. It records the highest scores on Clotho-AQA (70.11%) and TUT2017 (78.40%). On VocalSound, VideoLLaMA 2-AV (7B) reaches 93.19%, further narrowing the gap with Qwen2-Audio (93.92%). These results emphasize the effectiveness of the VideoLLaMA framework, showcasing both models’ ability to achieve superior or comparable results with much fewer training hours. This highlights their efficiency and demonstrates the value of integrating audio-visual modalities for enhanced learning.
@@
The performance results summarized in Table 8 highlight VideoLLaMA 2-AV (7B)’s significant advancements compared to open-source models. This analysis underscores VideoLLaMA 2-AV (7B)’s capabilities across various AVQA benchmarks, emphasizing its superior performance.
In the MUSIC-QA benchmark, VideoLLaMA 2-AV (7B) demonstrates substantial competency with a score of 79.2%. This score indicates a robust ability to analyze complex musical and audio cues, which is central to the MUSIC-QA benchmark. The model’s performance, outreaching CREAM with specialized model fine-tuning at 75.6%, underscores its advanced capabilities in handling intricate audio-visual interactions. In the AVSD benchmark, which focuses on audio-visual scene description involving dialogue understanding, VideoLLaMA 2-AV (7B) also demonstrates strong capabilities by scoring 57.2%, illustrating its adeptness at integrating visual context with audio inputs to generate coherent scene descriptions. The VGGSound benchmark, which necessitates a deep understanding of both visual scenes and their auditory elements, sees VideoLLaMA 2-AV (7B) leading with a score of 70.9%. This performance, much higher than AV-LLM’s 47.6%, showcases VideoLLaMA 2-AV (7B)’s exceptional ability to interpret and synthesize information across both modalities effectively. It indicates a profound understanding of dynamic interactions and the contextual synthesis necessary in AVSSD (audio-visual sound source detection) tasks.
@@ 7 Related Works
The vision-language adapters of Image-LLMs, such as Cross-Attention (Alayrac et al., 2022; Bai et al., 2023), Q-Former (Li et al., 2023a; Zhu et al., 2023b; Ye et al., 2024b), Linear Projection (Liu et al., 2023b; Chen et al., 2023a; Wang et al., 2023a) and Dynamic Visual Tokenizer (Jin et al., 2023c), have been widely adopted in popular Video-LLMs (Zhang et al., 2023; Maaz et al., 2023b; Lin et al., 2023a; Li et al., 2023c; Jin et al., 2024; Ataallah et al., 2024; Liu et al., 2024e). Despite their reasonable results, these designs are still insufficient for Video-LLMs because they completely ignore temporal aggregation and leave everything of temporal modeling to language decoder, which is neither computationally efficient (i.e., LLMs need to process much more tokens) nor effective on video understanding tasks, as observed in our preliminary experiments (See Table 1).
@@
Furthermore, expanding on the core capabilities of Video-LLMs, recent advancements have focused on integrating audio streams, a rich source of contextual cues that are vital for a complete video scene understanding. Models such as PandaGPT (Su et al., 2023) and Video-LLaMA (Zhang et al., 2023) utilize pre-trained systems like ImageBind (Girdhar et al., 2023) for a universal modality alignment, while MacawLLM (Lyu et al., 2023) optimizes separate branches for visual and audio data using diverse instructional sets. X-InstructBLIP (X-BLIP) (Panagopoulou et al., 2023) integrates multiple modalities within a single framework using frozen Large Language Models (LLMs), with modality-specific Q-Formers serving as adapters to bridge various encoders. OneLLM (Han et al., 2024) introduces a universal encoder and projection module designed to align multiple modalities with linguistic data, thereby improving the coherence of multimodal integration. CREMA (Yu et al., 2024) takes an efficient and modular approach, utilizing parameter-efficient adapters for each modality to enhance flexibility and ease of incorporating new modalities into existing frameworks. Other models like AV-LLM (Shu et al., 2023) and AVicuna (Tang et al., 2024) leverage integrated audio-visual training data to further refine their understanding of complex multimodal content. These developments represent significant strides toward creating more versatile and capable LLMs that can navigate and interpret intricate temporal dynamics with synchronous audio streams effectively.
@@ 8 Conclusion
We present the VideoLLaMA 2 series, a set of generalist Video Large Language Models designed to advance multimodal research in the arena of video-language modeling. By incorporating a meticulously designed Spatial-Temporal Convolution (STC) connector and a jointly trained Audio Branch, VideoLLaMA 2 consistently improves multimodal comprehension across various video and audio-oriented tasks. It outperforms the open-source models of similar size across multiple benchmarks and, in several aspects, achieves performance levels comparable to proprietary models, demonstrating impressive capabilities in modeling temporal dynamics with synchronous audio streams. Furthermore, as a foundational model, VideoLLaMA 2 can be further developed to benefit various more specialized but challenging problems, like long video understanding (Ren et al., 2023; Song et al., 2023; Wang et al., 2024d), video agent (Lin et al., 2023b; Fan et al., 2024; Wang et al., 2024b; He et al., 2024), autonomous driving (Xu et al., 2023b; Shao et al., 2024), motion understanding (Wu et al., 2024; Chen et al., 2024b), and robotic manipulation (Liu et al., 2024d).
@@ Appendix B Comparision Results with LLaVA-NeXT-Video Series
Appendix C Results on More Video Benchmarks
| Model | # Frames | MLVU | VideoVista | MMBench-Video | DREAM-1k |
| (M-AVG) | (Acc.) | (Score) | (F1) | ||
| VideoLLaMA 2 (7B) | 16 | 32.7 | 60.47 | 1.08 | 26.2 |
| VideoLLaMA 2 (72B) | 16 | 45.6 | - | - | 27.1 |